# Engineering morphogenesis of cell clusters with differentiable programming

### v2

@@ Abstract
Simulations run using the optimized model show a clear separation between the final costs of simulations run using untrained versus trained parameters (Supp Fig. S4b). Visualization of simulations confirms that trained parameters resulted in homogeneous growth (Fig. 5b). The learned gene circuit utilized both chemical and mechanical information to modulate division propensity (Fig. 5c, Supp Fig. S4d). To probe the robustness of the learned mechanism, we apply the trained gene circuit to simulations with other types of growth factor gradients: (a) an exponentially decaying growth factor gradient, (b) a gradient with two symmetric peaks, and (c) a gradient with a single asymmetric peak. Although we did not show any of these morphogen patterns during training, the gene circuit optimized on the power law gradient is still able to homogenize growth for the other types of gradients (Fig. 5d). Furthermore, the outputs of the gene network seem to track the shape of the fields, as well as patterns of stress accumulation (Fig. 5e, Supp Fig. S4c), consistently with a mechanical feedback mechanism. This is reminescent of the mechanism proposed by Hufnagel et al., where nonuniform growth leads to a buildup of mechanical stress at the center of the cluster, providing a feedback signal that cells can use to regulate their growth rates.
Chemical Regulation of Tissue Homeostasis
Maintaining stable ratios of various cell populations in tissue is important for proper functioning, and disruption of this can lead to disease [42]. To stably maintain homeostasis, the underlying cellular circuitry must prevent runaway growth for either cell type. Cells communicate via growth factors, and cell circuits that receive and send growth factors maintain tissue homeostasis, sometimes regenerating after perturbations. For example, co-cultures of macrophages and fibroblasts can reach stable ratios, even if initially plated with ratios that differ 100-fold [43]. Here, we aim at learning chemical mechanisms that can maintain a 50/50 ratio of two different cell types in a growing cell cluster.
In our model, two cell types are instantiated, each secreting one of two different chemicals. The simulation is seeded with an imbalanced initial state of $N_{init}=20$ total cells and $n_{r}=4$ red cells, and uniform division propensities. After 100 cell divisions, this results in a final cluster that preserves the initial imbalance (Fig 6a left). Gene interaction couplings are then optimized to minimize the number difference between the two cell types during growth (Fig 6a right). The loss at time $t$ takes the form:
$\mathcal{L}_{t}=|n_{r}(t)-n_{b}(t)|$
where $n_{r}$ and $n_{b}$ are the number of red and blue cells respectively, with $n_{r}+n_{b}=N$ being the total number of cells at time $t$.
[…]
Figure 6: Chemical Regulation of Homeostasis. (a) An initially imbalanced state (in the middle) grows into an equally unbalanced final state with untrained network parameters (left) and into a balanced state with learned parameters (right) (b) Visualization of the learned gene network. (c) Visualization of the chemical concentrations at the final state of the optimized simulation. (d) Average over 100 realizations of final proportion of red cells at the end of simulations of increasing lengths. (e) Average over 100 realizations of final proportion of red cells at the end of simulations with increasing proportions of red cells in the starting state. In both plots the shaded area corresponds to average plus or minus one standard deviation. (f) Visualization of learned gene networks when each cell type can learn its own network.
@@ Discussion
Our results demonstrate the power of gradient-based optimization methods in discovering mechanisms within the space of physical, cell-based models for collective behavior. By taking gradients through a physical model of cells interacting written using differentiable code, we were able to learn decision-making programs within individual cells to orchestrate growth into target shapes or with target properties. Additionally, we were able to overcome challenges associated with differentiating through the discrete nature of cell division events and non-differentiable loss functions through reinforcement learning based approaches. The mechanisms that were learned took the form of gene regulatory networks, which readily offer a biological interpretation. The learned mechanisms were local in nature; cells did not have access to global positional information and had to instead infer it by sensing chemicals and mechanics and communicating with nearby cells through these physical channels.
Although the gene networks optimized here are too large to instantiate experimentally, they could be distilled to find minimal networks or constrained to incorporate prior biological knowledge. Furthermore, the soft sphere model we have presented here is simply one of the possible choices. The physical details can be modified to optimize over an appropriate forward model with the same methods we demonstrated in this work. Questions about minimal mechanisms, robustness to variation and perturbations, number of unique solutions or growth principles could also be framed and investigated within the same setting.
Our results are a promising first step towards inferring mechanisms in existing multicellular systems or learning how to program them for desired behaviors. In particular, morphogenetic mechanisms frequently utilize local feedback to coordinate intricate forms, often through chemical or mechanical signals. Examples include a myriad of developmental systems, from the development of the wing imaginal disc in Drosophila to the periodic segmentation of blocks in somitogenesis. The development of robust and useful organoid systems also requires an understanding of spatiotemporal chemical and physical cues needed to guide the growth of stem cells. Physical models could be optimized to reproduce experimental data from time-lapse imaging of such developmental systems to gain insights into individual decision making functions that recapitulate complex behaviors. Gene expression datasets could be incorporated to bridge the gap between known and unknown information in biological systems. Our work opens up a multitude of avenues for further explorations of the landscape of growth programs for desired emergent behaviors in living systems.
@@
$\frac{dg_{i}}{dt}=\phi(\Sigma_{j}W_{ij}g_{j})+I_{i}-k_{i}g_{i}$
Considering the $i$-th gene, $g_{i}$ is the expression level, $W_{ij}$ is the regulatory influence of gene $j$ on $i$, $k_{i}$ is the degradation rate, and $I_{i}$ is a forcing signal coming from one of the sensed cell inputs. $\phi$ is a sigmoid function that ensures positive transcription rates saturating at high input levels.
@@
The gradient of simulations in Fig. 2 is calculated with standard automatic differentiation. The gradients of all the other simulations involve dealing with cell division events. Following REINFORCE [27], we can estimate the gradient of the expected loss for a single simulation:

### v3

@@ Abstract
The optimized gene circuit utilizes both chemical and mechanical information to modulate division propensity (Fig. 4b). This regulation leads to a significantly lower CV of cells’ division rates (Fig. 4c). Remarkably, the gene circuit optimized on the power law gradient successfully reduces growth variations also when exposed to external division scaling fields unseen during training. To assess the robustness of the mechanism, we expose the trained gene circuit to different external gradients: an exponentially decaying gradient, a gradient with two symmetric peaks, and a gradient with a single asymmetric peak.(Fig. 4c, left panel).
While simplified, our model parallels several phenomena observed in experimental studies of wing disc growth regulation and captures the essence of mechanical feedback in growth regulation. In the learned network, mechanical stress directly inhibits cell division, mirroring the role of the Hippo pathway in the wing disc. Experimentally, mechanical stress has been shown to activate the Hippo pathway [48], leading to Warts-mediated phosphorylation and inactivation of the growth-promoting transcription co-activator Yorkie (Yki), ultimately suppressing cell proliferation. Conversely, reduced mechanical stress allows Yki to remain unphosphorylated, enter the nucleus, and drive cell proliferation necessary for proper tissue growth and development. Our model represents a coarse-grained version of this mechanism through the inhibition of division propensity by the stress-sensing gene.
Although the secreted factors in our model lack direct biological counterparts, they serve as a simplified abstraction of the complex chemical signaling networks coordinating growth control in the Drosophila wing imaginal disc. The incorporation of secreted factors for feedback regulation conceptually represents the broader principle of growth control via morphogen gradients in the wing disc. For instance, the Decapentaplegic (Dpp) and Wingless (Wg) proteins form concentration gradients across the anterior-posterior and dorsal-ventral axes respectively, regulating growth and patterning [48]. While not direct analogues of these specific morphogens, our model’s diffusible factors capture the essence of spatially distributed chemical signals that influence growth patterns. The importance of these factors is further demonstrated in Fig. 4d, where we compare the performance of circuits with ablated inputs. Removing the chemical factors significantly impairs the cluster’s self-regulation capabilities, confirming that both chemical and mechanical regulation are essential in this context. Moreover, the gene network outputs appear to track both the shape of the fields and patterns of stress accumulation (Fig. 4e, Supp Fig. S6c), consistent with a mechanical feedback mechanism.
Additional Experiments
We report in the Supplementary Information a set of additional experimental setups that we omitted from the main text to focus on the most biologically relevant findings. We provide here a brief summary of such examples, to further stress the flexibility of our method and possibly inspire further applications. When cells have access to estimates of chemical gradients constructed by communicating with nearest neighbors, they can learn to elongate in one direction (Fig. S2) and grow in a branched shape (Fig. S3) by only including a single cell type. In these cases they learn to create emergent chemical gradients without the need for external asymmetric sources and show remarkable robustness to cell removal. We also show that, besides chemical secretion and division propensity, we can endow cells with control over their adhesion strength. By regulating a set of homotypic and heterotypic “cadherins”, cells can learn basic assembly programs for checkerboard, lobed and core-shell structures (Fig S1).
@@ Discussion
Our results demonstrate the power of gradient-based optimization methods in discovering mechanisms within the space of physical models for collective cell behavior. By taking gradients through a physical model of interacting cells written using automatically differentiable code, we were able to learn decision-making programs within individual cells to orchestrate growth into target shapes or with other target properties. We were able to overcome challenges associated with differentiating through the discrete nature of cell division events and non-differentiable loss functions with an approach based on classic reinforcement learning. The mechanisms that were learned took the form of gene regulatory networks, which readily offer a physical and biological interpretation. The learned mechanisms were local in nature: cells did not have access to global information (e.g. their absolute position) and had to infer it instead by sensing chemicals and mechanics and communicating with nearby cells through these physical channels.
The gene networks optimized here could be substituted with more detailed biophysical models or constrained to incorporate prior biological knowledge. Furthermore, the soft sphere model we have presented here is simply one of the possible choices. The physical details can be modified to optimize over an alternative bespoke forward model with the same methods we demonstrated in this work. Questions about minimal mechanisms, robustness to variation and perturbations, number of unique solutions or growth principles could also be framed and investigated within the same setting.
Our results are a promising first step towards inferring mechanisms in existing multicellular systems or learning how to program them for desired behaviors. In particular, morphogenetic mechanisms frequently utilize local feedback to coordinate intricate forms, often through chemical or mechanical signals. Examples include a myriad of developmental systems, from the development of the wing imaginal disc in Drosophila to the periodic segmentation of blocks in somitogenesis. The development of robust and useful organoid systems also requires an understanding of spatiotemporal chemical and physical cues needed to guide the growth of stem cells. Physical models could be optimized to reproduce experimental data from time-lapse imaging of such developmental systems to gain insights into individual decision making functions that recapitulate complex behaviors. Gene expression datasets could be incorporated to bridge the gap between known and unknown information in biological systems. Maximization or minimization principles could be explored beyond simplified models. Our work opens up a multitude of avenues for further explorations of the landscape of growth programs for desired emergent behaviors in living systems.
@@
$\frac{dg_{i}}{dt}=\phi(\Sigma_{j}W_{ij}g_{j}+b_{i})+I_{i}-k_{i}g_{i}$
Considering the $i$-th gene, $g_{i}$ is the expression level, $W_{ij}$ is the regulatory influence of gene $j$ on $i$, $b_{i}$ determines the gene transcription rate in absence of regulation, $k_{i}$ is the degradation rate, and $I_{i}$ is a forcing signal coming from one of the sensed cell inputs. $\phi$ is a sigmoid function that ensures positive transcription rates saturating at high input levels.
@@
Due to the presence of discrete stochastic events, the gradients of simulations involving cell division events cannot be calculated by direct application of the chain rule. Following REINFORCE [29], we can estimate the gradient of the expected loss for a single simulation: